OpenStudy (anonymous):

find the indicated probability a sample of 4 different calculators is randomly selected from a group containing 18 that are defective and 37 that have no defects. What is the probability that at least one of the calculators is defective?

OpenStudy (anonymous):

okay so i know i have to find P(x greaterthanorequalto 1)

OpenStudy (anonymous):

@jim_thompson5910 @Astrophysics @pooja195 @ganeshie8

jimthompson5910 (jim_thompson5910):

at least one? is that what you meant?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

do you agree that P(A) + P(B) = 1 where A = the event that at least one calculator is defective B = the event that no calculators are defective

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so we can say P(A) = 1 - P(B) we just need to figure out P(B)

OpenStudy (anonymous):

so the probability that no calculators are defective?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how many non-defective calcs are there?

OpenStudy (anonymous):

37 out of 55 which is 67% chance of one not being defective

jimthompson5910 (jim_thompson5910):

there are 37 non defective calcs how many ways are there to pick 4 of them (order doesn't matter) ?

OpenStudy (anonymous):

24?

jimthompson5910 (jim_thompson5910):

slot 1: 37 choices slot 2: 36 choices slot 3: 35 choices slot 4: 34 choices IF order mattered, then you'd have 37*36*35*34 = 1,585,080 different ways to select 4 of the 37 non defective calcs

jimthompson5910 (jim_thompson5910):

since order doesn't matter, you divide that by 4! = 24

OpenStudy (anonymous):

okay that's interesting what i did was multiply 4*3*2

OpenStudy (anonymous):

anyway what do i do now?

jimthompson5910 (jim_thompson5910):

what do you get when you divide by 24?

OpenStudy (anonymous):

do you mean divide the 67% ?

jimthompson5910 (jim_thompson5910):

divide 1,585,080 by 24

OpenStudy (anonymous):

okay 66045 , what does the 66045 represent?

jimthompson5910 (jim_thompson5910):

it represents the number of ways to pick 4 calculators out of the pool of 37

OpenStudy (anonymous):

okay okay so we now multiply this by the probability to get a non defective calc right? and then subtract that from 1?

jimthompson5910 (jim_thompson5910):

we do the same thing but now we include the 18 defective calcs so we have 18+37 = 55 total there are C(55,4) = 341,055 ways to pick 4 calcs (defective or nondefective) out of the pool of 55 The notation C(n,r) is the nCr combination notation

jimthompson5910 (jim_thompson5910):

B = the event that no calculators are defective P(B) = (# of ways to pick 4 non defective calculators)/(# of ways to pick any 4 calculators) P(B) = (66,045)/(341,055) P(B) = ???

OpenStudy (anonymous):

.19936

OpenStudy (anonymous):

and we subtract that from 1 right?

jimthompson5910 (jim_thompson5910):

I'm getting (66045)/(341055) = 0.19364911817742

OpenStudy (anonymous):

woops one too many nines , yes im getting that too

OpenStudy (anonymous):

and yep the answer is .806

OpenStudy (anonymous):

damn okay so how do you know what to do , i get that its logic and all but is their some special thought process that you go through to find the answer?

OpenStudy (anonymous):

how can i think more like you do?

jimthompson5910 (jim_thompson5910):

I think the first thing to determine is that you have to understand that P(A) + P(B) = 1 where A = the event that at least one calculator is defective B = the event that no calculators are defective events A and B are complementary events. Only one event must occur

jimthompson5910 (jim_thompson5910):

which leads to P(A) = 1 - P(B) to find P(B), you need to compute both C(37, 4) C(55, 4) You use the combination formula each time \[\Large C(n,r) = \frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

alright i need practice , thanks for you help really appreciate it :D

jimthompson5910 (jim_thompson5910):

no problem